# How smooth are restrictions of Besov functions?

### v1

@@ 3. Preliminary lemmata
$\Lambda_{\lfloor 2^{j}x\rfloor}^{*}=\frac{1}{\Psi(2^{-j})^{p}}\bigg{(}\sum_{k=1}^{j}\Psi(2^{-j})^{\varkappa}\bigg{)}^{L}\quad\text{for all }j\in J_{x}.$
@@
$\sup_{j\in\mathbb{N}}\,2^{\frac{j}{p}}\lambda_{j,\lfloor 2^{j}x\rfloor}\Psi(2^{-j})=\sup_{j\in\mathbb{N}}\,(\Lambda_{\lfloor 2^{j}x\rfloor}^{*})^{\frac{1}{p}}\Psi(2^{-j})\geqslant\sup_{j\in J_{x}}\bigg{(}\sum_{k=1}^{j}\Psi(2^{-j})^{\varkappa}\bigg{)}^{\frac{L}{p}}=\infty,$
@@
$\|f(\cdot,y)\|_{B_{p,\infty}^{(s,\Psi)}(\mathbb{R})}\gtrsim\sup_{j\in\mathbb{N}}\,2^{\frac{j}{p}}\lambda_{j,\lfloor 2^{j}y\rfloor}\Psi(2^{-j})=\infty\quad\text{for a.e. }y\in[1,2],$
so that $f(\cdot,y)\notin B_{p,\infty}^{(s,\Psi)}(\mathbb{R})$ for a.e. $y\in[1,2]$. (This uses only the representation of $B_{p,\infty}^{(s,\Psi)}(\mathbb{R})$ via differences, as in Definition 2.3.) This can be extended to a.e. $y\in\mathbb{R}$ (rather than “a.e. $y\in[1,2]$”) and to any dimension $d\in\mathbb{N}$ with $1\leqslant d<N$ using the same arguments as in [3].

### v3

@@ 3. Preliminary lemmata
$\Lambda_{\lfloor 2^{j}x\rfloor}^{*}=\frac{1}{\Psi(2^{-j})^{p}}\bigg(\sum_{k=1}^{j}\Psi(2^{-k})^{\varkappa}\bigg)^{L}\quad\text{for all }j\in J_{x}.$
@@
$\sup_{j\in\mathbb{N}}\,2^{\frac{j}{p}}\lambda_{j,\lfloor 2^{j}x\rfloor}\Psi(2^{-j})=\sup_{j\in\mathbb{N}}\,(\Lambda_{\lfloor 2^{j}x\rfloor}^{*})^{\frac{1}{p}}\Psi(2^{-j})\geqslant\sup_{j\in J_{x}}\bigg(\sum_{k=1}^{j}\Psi(2^{-k})^{\varkappa}\bigg)^{\frac{L}{p}}=\infty,$
@@
$\|f(\cdot,y)\|_{B_{p,\infty}^{(s,\Psi)}(\mathbb{R}^{N-1})}\gtrsim\sup_{j\in\mathbb{N}}\,2^{\frac{j}{p}}\lambda_{j,\lfloor 2^{j}y\rfloor}\Psi(2^{-j})=\infty\quad\text{for a.e. }y\in[1,2],$
so that $f(\cdot,y)\notin B_{p,\infty}^{(s,\Psi)}(\mathbb{R}^{N-1})$ for a.e. $y\in[1,2]$. (This uses only the representation of $B_{p,\infty}^{(s,\Psi)}(\mathbb{R}^{N-1})$ via differences, as in Definition 2.3.) This can be extended to a.e. $y\in\mathbb{R}$ (rather than “a.e. $y\in[1,2]$”) and to any dimension $d\in\mathbb{N}$ with $1\leqslant d<N$ using the same arguments as in [3].